OpenStudy (anonymous):

What is the value of f[g(3)] for the functions f(x) = x + 2 and g(x) = 5x - 1? (Type only the value of your final answer, no variables.) Numerical Answers Expected!

OpenStudy (anonymous):

f(g(3)) first, find g(3) . after that, put the result tumber on the formula. like this >> f(result g(3)) do u understand?

OpenStudy (anonymous):

not really.. can u help me work through this

OpenStudy (anonymous):

ok.. first.. find g(3) with this formula g(x) = 5x - 1

OpenStudy (anonymous):

soo what do i do now..

OpenStudy (anonymous):

find the result of g(3) just subtitute the x on this formula >> g(x) = 5x - 1 with 3. what is the result u got?

OpenStudy (anonymous):

so substitute with x with 3

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

g(3) = 5(3) - 1

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

so 14 is the answer ?

OpenStudy (anonymous):

it's the answer for g(3). now.. subtitute x with 14 to this formula f(x) = x + 2

OpenStudy (anonymous):

14+2 = 16

OpenStudy (anonymous):

so what now

OpenStudy (anonymous):

it's the final result

OpenStudy (anonymous):

so 16 is the answer not 14

OpenStudy (anonymous):

can u please help me with a different question

OpenStudy (anonymous):

f[g(3)] is f(g(3)) = g(3) + 2 1) to solve that first we must find g(3). g(3) = 5(3) - 1 g(3) = 15 - 1 = 14 2) after that, we can find f(g(3)). f(g(3)) = g(3) + 2 f(14) = 14 +2 f(14) = 16 do u understand?